First of all, here comes the original examples from the float package (just to see if they still work), the rotated versions of these will follow on the next pages (first without placement specifier, right afterwards with placement specifier [H]):

Program 0.1 The first program. This hasn’t got anything to do with the style but is included as an example. Note the ruled float style.

#include <stdio.h>
int main(int argc, char **argv)
{
       int i;
       for (i = 0; i < argc; ++i)
               printf("argv[%d] = %s\n", i, argv[i]);
       return 0;
}

     \floatstyle{ruled}
     \newfloat{Program}{tbp}{lop}[section]

… loads o’ stuff …

     \begin{Program}
     \begin{verbatim}

… program text …

     \end{verbatim}

\caption{… caption … }

     \end{Program}

Example 0.1: This is another silly floating Example.

\begin{matrix}n & \binom{n}{0} & \binom{n}{1} & \binom{n}{2} & \binom{n}{3} & \binom{n}{4} & \binom{n}{5} & \binom{n}{6} & \binom{n}{7} \\ 0 & 1 \\ 1 & 1 & 1 \\ 2 & 1 & 2 & 1 \\ 3 & 1 & 3 & 3 & 1 \\ 4 & 1 & 4 & 6 & 4 & 1 \\ 5 & 1 & 5 & 10 & 10 & 5 & 1 \\ 6 & 1 & 6 & 15 & 20 & 15 & 6 & 1 \\ 7 & 1 & 7 & 21 & 35 & 35 & 21 & 7 & 1 \end{matrix}

Table 1: Pascal’s triangle. This is a re-styled L^{A}TEX table.

Program 0.2 The first program. This hasn’t got anything to do with the style but is included as an example. Note the ruled float style.

#include <stdio.h>
int main(int argc, char **argv)
{
       int i;
       for (i = 0; i < argc; ++i)
               printf("argv[%d] = %s\n", i, argv[i]);
       return 0;
}

2

\floatstyle{ruled}
\newfloat{Program}{tbp}{lop}[section]

… loads o’ stuff …

\begin{Program}
\begin{verbatim}

… program text …

\end{verbatim}

\caption{… caption … }

\end{Program}

Example 0.2: This is another silly floating Example.

3

\begin{matrix}n & \binom{n}{0} & \binom{n}{1} & \binom{n}{2} & \binom{n}{3} & \binom{n}{4} & \binom{n}{5} & \binom{n}{6} & \binom{n}{7} \\ 0 & 1 \\ 1 & 1 & 1 \\ 2 & 1 & 2 & 1 \\ 3 & 1 & 3 & 3 & 1 \\ 4 & 1 & 4 & 6 & 4 & 1 \\ 5 & 1 & 5 & 10 & 10 & 5 & 1 \\ 6 & 1 & 6 & 15 & 20 & 15 & 6 & 1 \\ 7 & 1 & 7 & 21 & 35 & 35 & 21 & 7 & 1 \end{matrix}

Table 2: Pascal’s triangle. This is a re-styled L^{A}TEX table.

int main(int argc, char **argv)
{
       int i;
       for (i = 0; i < argc; ++i)
               printf("argv[%d] = %s\n", i, argv[i]);
       return 0;
}

5

\floatstyle{ruled}
\newfloat{Program}{tbp}{lop}[section]

… loads o’ stuff …

\begin{Program}
\begin{verbatim}

… program text …

\end{verbatim}

\caption{… caption … }

\end{Program}

Example 0.3: This is another silly floating Example.

6

\begin{matrix}n & \binom{n}{0} & \binom{n}{1} & \binom{n}{2} & \binom{n}{3} & \binom{n}{4} & \binom{n}{5} & \binom{n}{6} & \binom{n}{7} \\ 0 & 1 \\ 1 & 1 & 1 \\ 2 & 1 & 2 & 1 \\ 3 & 1 & 3 & 3 & 1 \\ 4 & 1 & 4 & 6 & 4 & 1 \\ 5 & 1 & 5 & 10 & 10 & 5 & 1 \\ 6 & 1 & 6 & 15 & 20 & 15 & 6 & 1 \\ 7 & 1 & 7 & 21 & 35 & 35 & 21 & 7 & 1 \end{matrix}

Table 3: Pascal’s triangle. This is a re-styled L^{A}TEX table.

THIS IS A FIGURE! 8 THIS IS A FIGURE! 9